# Pre-Generating Multi-Difficulty PDE Data for Few-Shot Neural PDE Solvers

### v1

@@ 1 Introduction
Deep learning has emerged as a powerful paradigm for solving PDEs, enabling data-driven surrogate models that can accelerate simulation, inference, and design across diverse scientific domains (li2021fourierneuraloperatorparametric; Lu_2021; pathak2022fourcastnetglobaldatadrivenhighresolution). This has been driven by the development of specialized models such neural operators (li2021fourierneuraloperatorparametric; Lu_2021) and recent transformer-based extensions (guibas2022adaptivefourierneuraloperators; brandstetter2023messagepassingneuralpde), which have demonstrated strong performance on benchmark datasets and gained traction in the machine learning for science community. More recently, significant effort has been devoted towards pre-generation of large datasets such as the Well (ohana2024well) and pre-training specialized foundation models (FMs) (herde2024poseidon; hao2024dpot; shen2024ups; mccabe2023multiple). The goal of these FMs is to serve as general-purpose foundations for PDE surrogates: delivering fast inference while minimizing—or even eliminating—the need to retrain on new, potentially harder-to-solve PDEs.
An underlying aspect of this line of work has long been the issue that it seeks to solve PDEs faster than classical numerical solvers but requires examples generated by the latter to do so. While such a circuitous setup is justifiable in many of the inverse problem applications that motivate learned solvers, it is still the case that tasks we eventually want to accelerate—practical engineering tasks in difficult-to-simulate regimes—will be exactly those for which it is hard to generate a significant amount. This need to decrease the sample complexity of neural PDE solvers has spurred significant research drawing up transfer learning (herde2024poseidon), active learning (bruna2024neural; musekamp2024active), and other method-centric approaches (rotman2023semi).
@@
Our first key result is that adding easy-to-medium difficulty data substantially improves performance on the hard distribution. Naturally, one might expect that medium difficulty data might be more useful, and our second main result is that there is often a favorable tradeoff justifying pre-generating medium-difficulty datasets instead of easy ones when solving cross in two classes of simulation, flow past an object (FPO) and lid-driven cavity (LDC), and using both supervised-only neural PDE solvers—specifically the Factorized Fourier Neural Operator (FFNO) (tran2021factorized) and the Convolutional Neural Operator (CNO) (raonic2023convolutional)—and the current state-of-the-art multi-physics pretrained Poseidon FM (herde2024poseidon). These complementary settings allow us to assess both specialized neural operators and large pretrained models under controlled difficulty-mixing regimes. In more detail, our contributions are the following:
@@
    Beyond square obstacles, we show that single simple‑obstacle data can improve the few-shot performance of models on flows around more complex shapes from FlowBench (tali2024flowbenchlargescalebenchmark), even when given only a few examples from it. This demonstrates the potential of a single dataset serving as a foundation for few-shot training of learned solvers on multiple harder datasets.
@@ 2.1 Difficulty axes
We next describe at a high level the settings used to generate the axes’ data; further details, including about our OpenFOAM (jasak2007openfoam) setup, can be found in Appendix B.
@@
Figure 2: FPO with objects from the FlowBench G1 NURBS data (tali2024flowbenchlargescalebenchmark).
@@
Lastly, we also enhance our investigation of geometry difficulty by changing the object complexity, specifically by incorporating simulations from FlowBench (tali2024flowbenchlargescalebenchmark). Specifically, we use their G1 dataset of FPO simulations of flows around Non-Uniform Rational B-Splines (NURBS), two examples of which are depicted in Figure 2. Because FlowBench is external, we do not measure the classical solve difficulty and treat no-obstacle and single (square) obstacle flows from the geometry axis as the easy and medium difficulty settings, respectively.
@@ 2.2 Pre-generated datasets
For each of the above settings and difficulty axes (excluding FlowBench) we generate $n=6,400$ simulations with different randomly sampled initial conditions. Specifically, following herde2024poseidon we sample Re between $[100,1000]$ (or higher if we are varying along the physics axis, as described above) and use that to set the inlet (FPO) or lid (LDC) velocity. We store the solution $\bm{y}_{t}^{i}=(\mathbf{u}_{i}(\mathbf{x},t),p_{i}(\mathbf{x},t))$ of each simulation $i=1,\dots,n$ at $T=20$ timesteps $t=1,\dots,T$ on a regular grid of points $\mathbf{x}\in\Omega$. Starting with this data, we hold out a subset of $N=100$ trajectories and set the goal of a learned PDE solver as using the remaining data to learn a model that, given the initial conditions $\bm{y}_{0}^{i}=(\mathbf{u}_{i}(\mathbf{x},0),p_{i}(\mathbf{x},0))$ of a held-out trajectory $i$, predicts a trajectory $\hat{\bm{y}}_{t}^{i}$ where $t\in[1,T]$ and $\hat{\bm{y}}_{t}^{i}=(\mathbf{\hat{u}}_{i}(\mathbf{x},t),\hat{p}_{i}(\mathbf{x},t))$. Following raonic2023convolutional; herde2024poseidon, we measure its success at doing so using the mean relative L1 error (nMAE):
@@ 3 Empirical results
The specific supervised models we consider are the Convolutional Neural Operator (CNO) (raonic2023convolutional) and the Factorized Fourier Neural Operator (FFNO) (tran2021factorized), which have demonstrated strong performance on several benchmarks (ohana2024well; tali2024flowbenchlargescalebenchmark; dauner2024resffno; takamoto2022pdebench; koehler2024apebench). These two models are trained from scratch on the different training mixtures we consider. To see whether our findings continue to hold in the higher performance regimes enabled by large-scale multi-physics pretraining, we also consider the Poseidon family of FMs trained on diverse PDE families (herde2024poseidon), evaluating three variants: Tiny (21M parameters), Base (158M), and Large (629M). Unlike CNO and FFNO, in this case we train or fine-tune the models on our training mixure starting from a model checkpoint pretrained on diverse PDE families. Training details of all models are reported in the Appendix.
@@ 3.3 Towards foundation datasets
In our last evaluation, we study the implications of multi-difficulty training for large-scale pre-generation of datasets for solving diverse PDE tasks. As an example of the latter, we consider examples from the FlowBench dataset (tali2024flowbenchlargescalebenchmark) of flows past irregular NURBS objects (cf. Section 2.1). Using flows past zero objects and flows past one square object as the easy and medium distributions, respectively, we show in Figure 11 that adding these simpler examples can reduce the error. This is especially pronounced in the case of FFNO when its training data is augmented with single object FPO examples. This suggests the potential utility of pre-generating large medium difficulty datasets and reusing them on multiple other datasets, thus amortizing the pre-generation costs. This pipeline is analogous to that of pre-training a foundation model, the cost of which is amortized as it is applied to multiple downstream tasks.
This idea of a pre-generated “foundation dataset” can also be used to describe much larger-scale efforts like The Well (ohana2024well). However, what our study demonstrates is that, just like the length and quality of web data used to train large language models matters, so does the difficulty of data pre-generated for training PDEs. In particular, as discussed in the last section, medium (across any axis) data can be much more effective as a mixing dataset than easy (across any axis) data for multi-obstacle performance. The current section further shows this for FlowBench NURBS data. Thus, when pre-generating such large datasets, it will be important to incorporate settings that more closely approach the types of more difficult problems that will be of actual interest to future users of neural PDE solvers.
@@ Appendix A Related work
Two leading paradigms for learning PDE solutions are physics-informed neural networks (PINNs) and neural operators. PINNs embed PDE residuals and boundary conditions into the loss, enabling mesh-free training, strong use of physics priors, and efficacy in small-data forward/inverse settings [raissi2019pinn, karniadakis2021]. At the same time, training can involve challenging multi-term loss balancing and optimization stiffness [krishnapriyan2021characterizing], with slower convergence on multi-scale or chaotic regimes (e.g., high Re turbulence) and sensitivity to complex geometries or boundary conditions [karniadakis2021]. Neural operators (e.g., DeepONet, FNO) learn mappings between function spaces, providing amortized inference, cross-discretization/geometry generalization, and scalability via pretraining on synthetic data [lu2021deeponet, li2021fno, kovachki2023]. Their performance, however, typically depends on substantial supervised datasets; robustness may be reduced under distribution shift across physics/geometry, and accuracy near shocks/discontinuities or conservation/stability guarantees may require additional structure and memory [kovachki2023, brandstetter2022, ruiz2024benefits]. Geometry-aware operator variants (e.g., GeoFNO) enhance robustness on irregular domains yet still rely on curated simulation corpora [li2023geofno]. Recent surveys synthesize these properties across PDE tasks, including turbulent flows [karniadakis2021, brunton2020].
@@
Unlike these efforts, we focus on the data itself, specifically on how generating better quality data may improve performance. There exists some work in the sciences on augmenting scarce experimental datasets with abundant simulated data from simplified systems, e.g. doi:10.1126/sciadv.aau6792 demonstrated that combining simulated flat-folding patterns with limited experimental data enabled machine learning models to recover structure in crumpled sheets. Similar studies are being conducted in biology, where recent work investigates the effects of training data composition on the performance of foundation models for single-cell genomics [nadig2025consequences]. We view our contribution as a more systematic study of how to generate and make use of data of different qualities.
@@
Outside of PDEs, data difficulty has also been explored in other areas such as language modeling. For example, a major difficulty axis in natural language processing is context length, with several explorations of how to train models capable of solving long-context tasks without resorting to purely long-context training [anil2022exploring, cho2024position]. Separately, our work is also related to work on easy-to-hard generalization in arithmetic reasoning tasks [sun2024easy, hase2024unreasonable] Here, the past work has found that appropriately training the models on simpler tasks—for example simpler math problems—leads to better performance on harder tasks. Here a task’s hardness is determined according to some human hardness measures, e.g. grade-level for STEM problems.
@@ B.1 FPO Domain
Figure 13: FPO Flow Fields (Velocity). Top row shows velocity and pressure fields for easy, medium, and complex setups; White areas represent the square obstacles (holes) in the domain.
@@ B.2 LDC Domain
Figure 14: LDC Flow Fields (Velocity). Top row: velocity and pressure fields for easy, medium, and complex setups.
@@ B.4.5 Design Motivation
This configuration follows established best practices in the OpenFOAM ecosystem and prior simulation-driven ML benchmarks, ensuring numerical stability and physical realism across a wide range of Reynolds numbers. We adopt Gauss linear for gradients and diffusive terms to preserve smoothness on structured grids weller1998tensor, and linearUpwind grad(U) to balance accuracy and robustness in the presence of sharp gradients and internal obstacles jasak1996error. The backward-Euler time integration and implicit solvers align with standard OpenFOAM settings for incompressible flows and are widely used in both industrial and academic studies jasak2007openfoam.
@@ Appendix C Model Architectures
In this section, we provide implementation and training details for the neural operator models evaluated in our study. We focus on two classes of models: (1) the Convolutional Neural Operator (CNO) [raonic2023convolutional], which is trained from scratch, and (2) Poseidon-T [herde2024poseidon], a pretrained transformer-based model fine-tuned on our downstream task. These models differ significantly in their architectural design, parameterization, and training strategy.
@@ C.1 Convolutional Neural Operator (CNO)
The Convolutional Neural Operator (CNO) used in our experiments is based on the architecture proposed by raonic2023convolutional, with several adjustments tailored to our time-dependent Navier-Stokes flows. The model is implemented in PyTorch Lightning and trained in an autoregressive supervised setting.
@@ C.2 Factorized Fourier Neural Operator (F-FNO)
The Factorized Fourier Neural Operator (F-FNO) used in our experiments is based on the architecture proposed by tran2021factorized, designed for efficient learned simulation of PDEs. The model is implemented in PyTorch Lightning and trained in an autoregressive one-step prediction setting. It consists of a deep sequence of Fourier operator layers with factorized spectral convolutions and improved residual connections, which allow stable training even at greater network depths than the original FNO . Key components of the architecture include:
@@
  • Coordinate Encoding: Following tran2021factorized, we augment the input with explicit spatial coordinate channels. Each input field is concatenated with its normalized $x$ and $y$ coordinates (as two additional channels), providing a positional encoding that consistently improves accuracy . The Fourier layers inherently utilize absolute positions (through the grid indices in the transform), while the feed-forward layers benefit from the coordinate features to capture location-dependent effects .
@@
  • Input Normalization and Noise: We apply per-variable normalization to input fields and add a small Gaussian noise perturbation during training . These techniques, recommended by tran2021factorized, act as regularization and help prevent training instabilities (we observed that without the added noise, the model’s validation loss could sometimes spike early in training ).

### v2

@@ 1 Introduction
Deep learning has emerged as a powerful paradigm for solving PDEs, enabling data-driven surrogate models that can accelerate simulation, inference, and design across diverse scientific domains (Li et al., 2021a; Lu et al., 2021a; Pathak et al., 2022). This has been driven by the development of specialized models such neural operators (Li et al., 2021a; Lu et al., 2021a) and recent transformer-based extensions (Guibas et al., 2022; Brandstetter et al., 2023), which have demonstrated strong performance on benchmark datasets and gained traction in the machine learning for science community. More recently, significant effort has been devoted towards pre-generation of large datasets such as the Well (Ohana et al., 2024) and pre-training specialized foundation models (FMs) (Herde et al., 2024; Hao et al., 2024; Shen et al., 2024; McCabe et al., 2023). The goal of these FMs is to serve as general-purpose foundations for PDE surrogates: delivering fast inference while minimizing—or even eliminating—the need to retrain on new, potentially harder-to-solve PDEs.
An underlying aspect of this line of work has long been the issue that it seeks to solve PDEs faster than classical numerical solvers but requires examples generated by the latter to do so. While such a circuitous setup is justifiable in many of the inverse problem applications that motivate learned solvers, it is still the case that tasks we eventually want to accelerate—practical engineering tasks in difficult-to-simulate regimes—will be exactly those for which it is hard to generate a significant amount. This need to decrease the sample complexity of neural PDE solvers has spurred significant research drawing up transfer learning (Herde et al., 2024), active learning (Bruna et al., 2024; Musekamp et al., 2024), and other method-centric approaches (Rotman et al., 2023).
@@
Our first key result is that adding easy-to-medium difficulty data substantially improves performance on the hard distribution. Naturally, one might expect that medium difficulty data might be more useful, and our second main result is that there is often a favorable tradeoff justifying pre-generating medium-difficulty datasets instead of easy ones when solving cross in two classes of simulation, flow past an object (FPO) and lid-driven cavity (LDC), and using both supervised-only neural PDE solvers—specifically the Factorized Fourier Neural Operator (FFNO) (Tran et al., 2021) and the Convolutional Neural Operator (CNO) (Raonic et al., 2023)—and the current state-of-the-art multi-physics pretrained Poseidon FM (Herde et al., 2024). These complementary settings allow us to assess both specialized neural operators and large pretrained models under controlled difficulty-mixing regimes. In more detail, our contributions are the following:
@@
    Beyond square obstacles, we show that single simple‑obstacle data can improve the few-shot performance of models on flows around more complex shapes from FlowBench (Tali et al., 2024), even when given only a few examples from it. This demonstrates the potential of a single dataset serving as a foundation for few-shot training of learned solvers on multiple harder datasets.
@@ 2.1 Difficulty axes
We next describe at a high level the settings used to generate the axes’ data; further details, including about our OpenFOAM (Jasak et al., 2007) setup, can be found in Appendix B.
@@
Figure 2: FPO with objects from the FlowBench G1 NURBS data (Tali et al., 2024).
@@
Lastly, we also enhance our investigation of geometry difficulty by changing the object complexity, specifically by incorporating simulations from FlowBench (Tali et al., 2024). Specifically, we use their G1 dataset of FPO simulations of flows around Non-Uniform Rational B-Splines (NURBS), two examples of which are depicted in Figure 2. Because FlowBench is external, we do not measure the classical solve difficulty and treat no-obstacle and single (square) obstacle flows from the geometry axis as the easy and medium difficulty settings, respectively.
@@ 2.2 Pre-generated datasets
For each of the above settings and difficulty axes (excluding FlowBench) we generate $n=6,400$ simulations with different randomly sampled initial conditions. Specifically, following Herde et al. (2024) we sample Re between $[100,1000]$ (or higher if we are varying along the physics axis, as described above) and use that to set the inlet (FPO) or lid (LDC) velocity. We store the solution $\bm{y}_{t}^{i}=(\mathbf{u}_{i}(\mathbf{x},t),p_{i}(\mathbf{x},t))$ of each simulation $i=1,\dots,n$ at $T=20$ timesteps $t=1,\dots,T$ on a regular grid of points $\mathbf{x}\in\Omega$. Starting with this data, we hold out a subset of $N=100$ trajectories and set the goal of a learned PDE solver as using the remaining data to learn a model that, given the initial conditions $\bm{y}_{0}^{i}=(\mathbf{u}_{i}(\mathbf{x},0),p_{i}(\mathbf{x},0))$ of a held-out trajectory $i$, predicts a trajectory $\hat{\bm{y}}_{t}^{i}$ where $t\in[1,T]$ and $\hat{\bm{y}}_{t}^{i}=(\mathbf{\hat{u}}_{i}(\mathbf{x},t),\hat{p}_{i}(\mathbf{x},t))$. Following Raonic et al. (2023); Herde et al. (2024), we measure its success at doing so using the mean relative L1 error (nMAE):
@@ 3 Empirical results
The specific supervised models we consider are the Convolutional Neural Operator (CNO) (Raonic et al., 2023) and the Factorized Fourier Neural Operator (FFNO) (Tran et al., 2021), which have demonstrated strong performance on several benchmarks (Ohana et al., 2024; Tali et al., 2024; Dauner and others, 2024; Takamoto et al., 2022; Koehler and others, 2024). These two models are trained from scratch on the different training mixtures we consider. To see whether our findings continue to hold in the higher performance regimes enabled by large-scale multi-physics pretraining, we also consider the Poseidon family of FMs trained on diverse PDE families (Herde et al., 2024), evaluating three variants: Tiny (21M parameters), Base (158M), and Large (629M). Unlike CNO and FFNO, in this case we train or fine-tune the models on our training mixure starting from a model checkpoint pretrained on diverse PDE families. Training details of all models are reported in the Appendix.
@@ 3.3 Towards foundation datasets
In our last evaluation, we study the implications of multi-difficulty training for large-scale pre-generation of datasets for solving diverse PDE tasks. As an example of the latter, we consider examples from the FlowBench dataset (Tali et al., 2024) of flows past irregular NURBS objects (cf. Section 2.1). Using flows past zero objects and flows past one square object as the easy and medium distributions, respectively, we show in Figure 11 that adding these simpler examples can reduce the error. This is especially pronounced in the case of FFNO when its training data is augmented with single object FPO examples. This suggests the potential utility of pre-generating large medium difficulty datasets and reusing them on multiple other datasets, thus amortizing the pre-generation costs. This pipeline is analogous to that of pre-training a foundation model, the cost of which is amortized as it is applied to multiple downstream tasks.
This idea of a pre-generated “foundation dataset” can also be used to describe much larger-scale efforts like The Well (Ohana et al., 2024). However, what our study demonstrates is that, just like the length and quality of web data used to train large language models matters, so does the difficulty of data pre-generated for training PDEs. In particular, as discussed in the last section, medium (across any axis) data can be much more effective as a mixing dataset than easy (across any axis) data for multi-obstacle performance. The current section further shows this for FlowBench NURBS data. Thus, when pre-generating such large datasets, it will be important to incorporate settings that more closely approach the types of more difficult problems that will be of actual interest to future users of neural PDE solvers.
@@ Appendix A Related work
Two leading paradigms for learning PDE solutions are physics-informed neural networks (PINNs) and neural operators. PINNs embed PDE residuals and boundary conditions into the loss, enabling mesh-free training, strong use of physics priors, and efficacy in small-data forward/inverse settings [Raissi et al., 2019, Karniadakis et al., 2021]. At the same time, training can involve challenging multi-term loss balancing and optimization stiffness [Krishnapriyan et al., 2021], with slower convergence on multi-scale or chaotic regimes (e.g., high Re turbulence) and sensitivity to complex geometries or boundary conditions [Karniadakis et al., 2021]. Neural operators (e.g., DeepONet, FNO) learn mappings between function spaces, providing amortized inference, cross-discretization/geometry generalization, and scalability via pretraining on synthetic data [Lu et al., 2021b, Li et al., 2021b, Kovachki et al., 2023]. Their performance, however, typically depends on substantial supervised datasets; robustness may be reduced under distribution shift across physics/geometry, and accuracy near shocks/discontinuities or conservation/stability guarantees may require additional structure and memory [Kovachki et al., 2023, Brandstetter et al., 2022, Ruiz et al., 2024]. Geometry-aware operator variants (e.g., GeoFNO) enhance robustness on irregular domains yet still rely on curated simulation corpora [Li et al., 2023]. Recent surveys synthesize these properties across PDE tasks, including turbulent flows [Karniadakis et al., 2021, Brunton and Kutz, 2020].
@@
Unlike these efforts, we focus on the data itself, specifically on how generating better quality data may improve performance. There exists some work in the sciences on augmenting scarce experimental datasets with abundant simulated data from simplified systems, e.g. Hoffmann et al. [2019] demonstrated that combining simulated flat-folding patterns with limited experimental data enabled machine learning models to recover structure in crumpled sheets. Similar studies are being conducted in biology, where recent work investigates the effects of training data composition on the performance of foundation models for single-cell genomics [Nadig et al., 2025]. We view our contribution as a more systematic study of how to generate and make use of data of different qualities.
@@
Outside of PDEs, data difficulty has also been explored in other areas such as language modeling. For example, a major difficulty axis in natural language processing is context length, with several explorations of how to train models capable of solving long-context tasks without resorting to purely long-context training [Anil et al., 2022, Cho et al., 2024]. Separately, our work is also related to work on easy-to-hard generalization in arithmetic reasoning tasks [Sun et al., 2024, Hase et al., 2024] Here, the past work has found that appropriately training the models on simpler tasks—for example simpler math problems—leads to better performance on harder tasks. Here a task’s hardness is determined according to some human hardness measures, e.g. grade-level for STEM problems.
@@ B.1 FPO Domain
Figure 13: FPO Flow Fields (Velocity). Top row shows velocity and relative pressure fields for easy, medium, and complex setups; White areas represent the square obstacles (holes) in the domain. Note: The visualized pressure represents relative (gauge) pressure and may be negative, as the incompressible Navier-Stokes equations depend only on pressure gradients rather than absolute pressure.
@@ B.2 LDC Domain
Figure 14: LDC Flow Fields (Velocity). Top row: velocity and relative pressure fields for easy, medium, and complex setups. Note that pressure is shown as relative (gauge) pressure, which allows for negative values consistent with incompressible flow formulation.
@@ B.4.5 Design Motivation
This configuration follows established best practices in the OpenFOAM ecosystem and prior simulation-driven ML benchmarks, ensuring numerical stability and physical realism across a wide range of Reynolds numbers. We adopt Gauss linear for gradients and diffusive terms to preserve smoothness on structured grids Weller et al. [1998], and linearUpwind grad(U) to balance accuracy and robustness in the presence of sharp gradients and internal obstacles Jasak [1996]. The backward-Euler time integration and implicit solvers align with standard OpenFOAM settings for incompressible flows and are widely used in both industrial and academic studies Jasak et al. [2007].
@@ Appendix C Model Architectures
In this section, we provide implementation and training details for the neural operator models evaluated in our study. We focus on two classes of models: (1) the Convolutional Neural Operator (CNO) [Raonic et al., 2023], which is trained from scratch, and (2) Poseidon-T [Herde et al., 2024], a pretrained transformer-based model fine-tuned on our downstream task. These models differ significantly in their architectural design, parameterization, and training strategy.
@@ C.1 Convolutional Neural Operator (CNO)
The Convolutional Neural Operator (CNO) used in our experiments is based on the architecture proposed by Raonic et al. [2023], with several adjustments tailored to our time-dependent Navier-Stokes flows. The model is implemented in PyTorch Lightning and trained in an autoregressive supervised setting.
@@ C.2 Factorized Fourier Neural Operator (F-FNO)
The Factorized Fourier Neural Operator (F-FNO) used in our experiments is based on the architecture proposed by Tran et al. [2021], designed for efficient learned simulation of PDEs. The model is implemented in PyTorch Lightning and trained in an autoregressive one-step prediction setting. It consists of a deep sequence of Fourier operator layers with factorized spectral convolutions and improved residual connections, which allow stable training even at greater network depths than the original FNO . Key components of the architecture include:
@@
  • Coordinate Encoding: Following Tran et al. [2021], we augment the input with explicit spatial coordinate channels. Each input field is concatenated with its normalized $x$ and $y$ coordinates (as two additional channels), providing a positional encoding that consistently improves accuracy . The Fourier layers inherently utilize absolute positions (through the grid indices in the transform), while the feed-forward layers benefit from the coordinate features to capture location-dependent effects .
@@
  • Input Normalization and Noise: We apply per-variable normalization to input fields and add a small Gaussian noise perturbation during training . These techniques, recommended by Tran et al. [2021], act as regularization and help prevent training instabilities (we observed that without the added noise, the model’s validation loss could sometimes spike early in training ).